OpenStudy (gahm8684):

If the quantity 4 times x times y cubed plus 8 times x squared times y to the fifth power all over 2 times x times y squared is completely simplified to 2xayb + 4xcyd, where a, b, c, and d represent integer exponents, what is the value of a? _______ Numerical Answers Expected

OpenStudy (gahm8684):

\[\frac{ 4yx^3+8x^2y^5 }{ 2xy^2 }\]

OpenStudy (gahm8684):

is actually 4xy^2

OpenStudy (gahm8684):

@anyone:c

OpenStudy (gahm8684):

@quickstudent @windsylph @esshotwired @radar @tejasvir @inkyvoyd @ObeySoph▲ @phi @Ashleyisakitty @SolomonZelman @sammixboo @dan815 @Gin_Ichimaru @Hero @jdoe0001 @kropot72 @lopezking1 @ztp5031 @xoxojac @Compassionate @VortexAlliby @bibby @nincompoop @Mchilds15

OpenStudy (gahm8684):

help please!!:c and sorry to bother you guys

OpenStudy (quickstudent):

Sorry, I'm not sure how to do this.

OpenStudy (gahm8684):

its ok

OpenStudy (quickstudent):

@campbell_st ?

OpenStudy (gahm8684):

@superhelp101 @nikato @lankymonkey

OpenStudy (superhelp101):

Value of a is zero

OpenStudy (anonymous):

agrees with superhelp \[\large \frac{ 2xy^2 (2y+4xy^3) }{ 2xy^2 }\] cancels the 2xy^2 on top and bottom \[\large =2y+4xy^3= 2x^0y^1 + 4x^1y^3\] so x^a= x^0 a=0

OpenStudy (anonymous):

@superhelp101 the drawing! hold me for a sec >.< <333

OpenStudy (anonymous):

I mean hold on*

OpenStudy (superhelp101):

Yayay nice job bestie!!!!

OpenStudy (anonymous):

you too bestie! <333 gimme a sec to take pic and upload! >.<